OpenStudy (ray1998):

Can someone please help me with this Algebraic question? I need to be walked through it, because I have no clue how to solve for it. This is for Algebra 2. 1. Solve the system by Elimination. x + 5y - 4z = -10 2x - y + 5z = -9 2x - 10y - 5z = 0 Thank you in advance and whomever is able to help will receive medal :)

OpenStudy (civicsiscool44):

yes

OpenStudy (ray1998):

Okay, so what do we do? @civicsiscool44

OpenStudy (civicsiscool44):

what are your answer choices if any

OpenStudy (ray1998):

Okay! Hold on.

OpenStudy (ray1998):

A. (5, -1, 0) B. (-5, 1, 0) C. (-5, -1, 0) D. (-5, -1, -2)

OpenStudy (ray1998):

@civicsiscool44

OpenStudy (civicsiscool44):

I'M 78% sure its d

OpenStudy (civicsiscool44):

don't forget to fan me

OpenStudy (ray1998):

Okay. Do you think you could explain a little to me how to do it? @civicsiscool44

OpenStudy (civicsiscool44):

well i have done this before and i know the answer is right

OpenStudy (ray1998):

It wasn't that I didn't trust your judgement, just that I really need help with knowing how to do this stuff lol xD but yeh, thank you. I'll give you your medal for trying. :) thank you! @civicsiscool44

OpenStudy (ray1998):

@ospreytriple , do you think you could help explain the problem to me? Like, how to solve it?

OpenStudy (civicsiscool44):

not really i'm really good at explaining math problems

OpenStudy (anonymous):

The method of elimination involves adding or subtracting entire equations in order to eliminate one of the variables. If I may refer to the equations in your problem as equations 1, 2, and 3, you can add equations 2 & 3 together to eliminate the variable z.

OpenStudy (ray1998):

Okay. Okay, gotcha @ospreytriple

OpenStudy (anonymous):

2x - y + 5z = -9 2x - 10y - 5z = 0 ------------------ 4x - 11y = -9 Get the idea?

OpenStudy (ray1998):

Yeh, I'm starting to a little. Thank you :) @ospreytriple

OpenStudy (anonymous):

Next, you'll have to add or subtract a different combination of the three equations to eliminate z from. Can you do that?

OpenStudy (anonymous):

How about multiplying equation 1 by 5 and equation 2 by 4 and then adding the two results together?

OpenStudy (ray1998):

I'm trying to think of how I would do that, but I'm not sure. I don't know where to start. @ospreytriple

OpenStudy (anonymous):

OK. Let's work with equations 1 and 2 and try to eliminate the z term.

OpenStudy (anonymous):

Notice that in equation 1, the coefficient of z is 4 and in equation 2, the coefficient of z is 5. We can't eliminate z unless the coefficient is the same in both equations.

OpenStudy (ray1998):

Okay. So do we move on to the ys and xs?

OpenStudy (anonymous):

Not yet. We have to eliminate z again from a different pair of equations. As I'm trying to explain, we can attempt to eliminate z by combining multiples of equations 1 and 2. Understand?

OpenStudy (ray1998):

Ohh yes. Equations 1 and 2, both share 5z.

OpenStudy (anonymous):

No. Equation 1 has 4z and equation 2 has 5 z. We have to make the coefficients the same.

OpenStudy (ray1998):

Not 1 and 2, but 2 and 3, I mean.

OpenStudy (ray1998):

Ohh, okay.

OpenStudy (anonymous):

The LCM of 4 and 5 is 20, so multiply equation 1 by 5 to make its z coefficient 20. Then multiply equation 2 by 5 to make its z coefficient 20. Can you do that?

OpenStudy (anonymous):

Sirry. Multiply equation 2 by 4.

OpenStudy (anonymous):

What do you get if you multiply equation 1 by 5?

OpenStudy (ray1998):

20?

OpenStudy (anonymous):

No. Multiply the whole equation...all the terms. x(5) + 5y(5) - 4z(5) = -10(5) What does that give you?

OpenStudy (ray1998):

1 + 25 - 20 = -50? But that wouldn't work, because of the different variables..

OpenStudy (ray1998):

5* + 25 - 20 = -50

OpenStudy (anonymous):

No. You can't drop the variables. For example, what do you get when you multiply\[3x \times 4 = ?\]

OpenStudy (ray1998):

12x ?

OpenStudy (anonymous):

Very good. Now try multiplying the whole equation 1 by 5. You got the coefficients correct a moment ago, but you can't drop the variables.

OpenStudy (ray1998):

5x + 25y - 20z = -50

OpenStudy (anonymous):

Perfect. Maybe call this equation 4. Now, in exactly the same way, multiply equation 3 by 4. What do you get?

OpenStudy (anonymous):

Sorry. Oops. Multiply equation 2 by 4. My bad.

OpenStudy (ray1998):

Oh. Okay lol you're good, thank you for being patient with me. 8x - 4y + 20z?

OpenStudy (anonymous):

Excellent. But don't forget the -36 on the right hand side. Call this equation 5. Carefully examine equations 4 & 5. If you add them together, you will eliminate the z term again. Can you do that?

OpenStudy (ray1998):

Oh, we include the = 9 part? Gotcha. Okay, hold on.

OpenStudy (ray1998):

13x + 29y + 40z? Or do we take both the 20s and cross them out?

OpenStudy (anonymous):

Add these two: 5x + 25y - 20z = -50 8x - 4y + 20z = -36 --------------------

OpenStudy (ray1998):

13x - 29y + 40z = 86

OpenStudy (anonymous):

Not quite. What do you get when you add 25 and -4 ?

OpenStudy (ray1998):

Opps, sorry! 21y

OpenStudy (anonymous):

Good. Now add 20z and -20z. What do you get?

OpenStudy (ray1998):

That's zero

OpenStudy (anonymous):

Yayyy! ELIMINATION. Now, on the right hand side, add -50 and -36. What do you get?

OpenStudy (ray1998):

Ahaha ^.^ wait, so you add -36 to each side?

OpenStudy (ray1998):

so, -50 -36 -36 -36 ?

OpenStudy (anonymous):

NO! Add the right hand sides together just like all the other terms. What do you get when you add -50 and -36?

OpenStudy (ray1998):

86

OpenStudy (anonymous):

How about -86?

OpenStudy (ray1998):

Oh okay, yes

OpenStudy (anonymous):

Alright. To sum up, the addition of equations 4 & 5 looks like this: 5x + 25y - 20z = -50 8x - 4y + 20z = -36 -------------------- 13x + 21y = -86 You OK with this? We're getting closer to the answer.

OpenStudy (ray1998):

Yes. Were good :) sorry for my slowness lol.

OpenStudy (anonymous):

No problem. Stay with me for another few minutes and we'll be done. What we've done so far: Added equations 2 & 3 together to give 4x - 11y = -9 Added 5 times equation 1 and 4 times equation 2 to give 13x + 21y = -86 Now we have completely eliminated z from the problem. We're down to only two equations with only x & y in them. Let's now try to eliminate y, OK?

OpenStudy (ray1998):

Okay

OpenStudy (anonymous):

Notice that the first equation above has -11y and the second equation above has +21y. In order to eliminate y we have to make the coefficients the same. The LCM of 11 and 21 is 231. So you'll have to multiply the first equation by 21 and the second equation by 11. Can you do that?

OpenStudy (ray1998):

Yeh, hold on. Let me try and work it out on paper :)

OpenStudy (anonymous):

Take your time...

OpenStudy (ray1998):

84x -231y = -189 143x + 231y = -946

OpenStudy (anonymous):

Excellent. Now add these two equations together to eliminate y. What do you get?

OpenStudy (ray1998):

Okay. 84x -231y = -189 143x + 231y = -946 227x + 0 = -757

OpenStudy (anonymous):

Check the right hand side. Add -189 and -946.

OpenStudy (ray1998):

-179572

OpenStudy (anonymous):

No. Are you using a calculator?

OpenStudy (ray1998):

Yes

OpenStudy (anonymous):

OK try again. -189 + -946 = ?

OpenStudy (ray1998):

1135?

OpenStudy (anonymous):

How about -1135?

OpenStudy (ray1998):

That's what I got doing it on paper, cause my calculator kept giving me different answers

OpenStudy (ray1998):

Yes. Sorry, I keep thinking that two negatives added together make a positive..

OpenStudy (anonymous):

Good. So now you have 227x = -1135. y and z have been eliminated. Can you solve for x?

OpenStudy (ray1998):

227x -1135 227x 227x 227x crosses our and becomes zero. x = 1130?

OpenStudy (anonymous):

No, to solve\[227x = -1135\]you would divide both sides by 227, i.e.\[\frac{ 227x }{ 227 } = \frac{ -1135 }{ 227 }\]\[x = ?\]

OpenStudy (ray1998):

That's what I meant.

OpenStudy (anonymous):

What do you get when you divide -1135 by 227 ?

OpenStudy (ray1998):

Sorry, it's my stupid calculator. I think it's broken. Hold on :) -5?

OpenStudy (anonymous):

That's right, x = -5. Now go back and find one of your equations with just x and y in it. Put -5 in for x and solve for y. Pick the simplest equation you can find.

OpenStudy (ray1998):

Just a random equation, and do this? 4x(-5) - 11y = -9

OpenStudy (anonymous):

That's right, but there won't be any x. You'll replace x with -5.

OpenStudy (anonymous):

4(-5) - 11y = -9 Solve for y.

OpenStudy (ray1998):

Ohhh, okay! Gotcha. Thanks :)

OpenStudy (anonymous):

What do you get for y?

OpenStudy (ray1998):

-20 - 11y = -9 11y 11y -20 - 0 = ?

OpenStudy (ray1998):

Do I multiply or divide?

OpenStudy (anonymous):

-20 - 11y = -9 First, add 20 to both sides. What do you get?

OpenStudy (ray1998):

-20 - 11y = -9 -20 -20 20 crosses out and becomes 0. Then, 11y + -20 = -9

OpenStudy (anonymous):

No. You added 20 twice to the left hand side. Just add 20 to left hand side and add 20 to the right had side, -20 - 11y + 20 = -9 + 20 What do you end up with?

OpenStudy (ray1998):

31 + 11

OpenStudy (anonymous):

No. Look at just the left hand side. You have a -20 and a +20. When you add them, what do you get?

OpenStudy (ray1998):

0